OpenStudy (anonymous):

helps medal and fan Match each item in Column A to an answer in Column B. What property is shown by the equation? column A --abc 1. (34 + 11) + 9 = (11 + 34) + 9 --abc 2. 25 + (-15 + 4) = (25 + (-15)) + 4 --abc 3. -(-21 + 8) = 21 + (-8) --abc 4. -30 + (-16) = -16 + (-30) column B A. commutative property B. associative property C. opposite of a sum property

OpenStudy (anonymous):

\[a + (b+c) = (a + b) + c \quad \quad \text{Associative Property}\]

OpenStudy (anonymous):

1 is b

OpenStudy (anonymous):

??????????

OpenStudy (anonymous):

Yes, you are right..

OpenStudy (anonymous):

What about 2??

OpenStudy (anonymous):

c???

OpenStudy (anonymous):

2 is c ??

OpenStudy (anonymous):

Are you seeing any difference between 1 and 2? Just there right hand side is on left hand side and left on right with respect to 1, right??

OpenStudy (anonymous):

the parinthisis change numbers

OpenStudy (anonymous):

See, whether \(a + (b+c) = (a+b) + c\) or \((a+b) + c = a+(b+c)\), they both are one and the same, and they come under Associative Property..

OpenStudy (anonymous):

ohh so 2 is b

OpenStudy (anonymous):

Parenthesis are only changed there, it is same as first case only..

OpenStudy (anonymous):

what is communitive property and oppisit of a sum property

OpenStudy (anonymous):

I don't know about opposite of sum property, and Commutative Property says: \[a+ b = b+a\]

OpenStudy (anonymous):

so 3 is a

OpenStudy (anonymous):

Actually not...

OpenStudy (anonymous):

no wait 4 is b meaning 3 is c

OpenStudy (anonymous):

You can say it for 4th one...

OpenStudy (anonymous):

mess up i feel stupid i ment 4 is a and 3 is c

OpenStudy (anonymous):

Yes, for 3 you can say that 3 is opposite of Sum Property.. May be multiply by -1 is Opposite of Sum Property ie: \[-(a+b) = -a -c\]

OpenStudy (anonymous):

Yes, 4 is a and 3 is c.. Good.. 1 and 2 are b.. Cleared?

OpenStudy (anonymous):

tanx

OpenStudy (anonymous):

Actually to my knowledge, multiply comes under Distribution.. Distribution Property says: \[a(b+c) = ab + ac \quad \quad \text{a= -1 \in your question}\]

OpenStudy (anonymous):

Sorry, the property name is : Distributive Property not Distribution.. :) Welcome.. :)

OpenStudy (anonymous):

What is the value of the expression? 12.4 + (–2 + 1.6)

OpenStudy (anonymous):

What is the value of the expression? 12.4 + (-2 + 1.6)